OpenStudy (anonymous):

what is 0.6 repeating as a percent?

OpenStudy (anonymous):

0.006 repeating.

OpenStudy (anonymous):

im confused ?

OpenStudy (anonymous):

66.6 % repeating

OpenStudy (anonymous):

If u don't like that one, then 0. 6 repeating is 2/3 but you have already been explained this by both Amistre and me.

OpenStudy (anonymous):

66.6 repeating %

OpenStudy (anonymous):

yeah i know and i dident understand what u were trying to tell me with that other problem!

OpenStudy (anonymous):

If I say 1 want 1% of something, I want 1/100 of it.

OpenStudy (anonymous):

0.5 expressed in percents is 50% isn't it ?

OpenStudy (amistre64):

the key to understanding is in responding with questions :)

OpenStudy (anonymous):

0.5 = 50/100

OpenStudy (anonymous):

0.5=50/100=50%

OpenStudy (anonymous):

No questions...

OpenStudy (anonymous):

So 0.66666666 (repeating) expressed in percents = 66.66 (repeating) % right?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

how do u put a repeating sign over a number?

OpenStudy (anonymous):

on the computer

OpenStudy (anonymous):

\[\bar 3\]

OpenStudy (anonymous):

like that ^ :)

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

type \bar 3 in the equation editor or put it inside \[\] and type it directly.

OpenStudy (anonymous):

It won't let me type the symbols, to put it direcly u put it between backslash open square bracket and backslash close square bracket.

OpenStudy (anonymous):

Do u know this site? http://www.purplemath.com/modules/percents.htm

OpenStudy (anonymous):

umm no but i can go to it